OpenStudy (anonymous):

first rewrite 14/15 and 9/10 so that they have common denominator then use <,= or > to order 14/15 and 9/10

OpenStudy (anonymous):

\[\frac{ 14 }{15 } \times \frac{ 10 }{ 10 } and \frac{ 9 }{ 10 } \times \frac{ 15 }{ 15 }\] \[\frac{ 140 }{ 150 } > \frac{ 135 }{ 150 }\]

OpenStudy (anonymous):

thanks